OpenStudy (karlaltr):

The motor of an elevator can produce 2200w of power. The elevator has a mass of 1100kg complete with contents. At what constant speed will the elevator rise?

OpenStudy (anonymous):

Do they give you an equation?

OpenStudy (karlaltr):

no:/

OpenStudy (dan815):

power is the rate of work

OpenStudy (dan815):

The unit for watts is Jules/sec = N*m/sec

OpenStudy (dan815):

can u figure it out now?

OpenStudy (karlaltr):

im confused because wouldnt we be using a formula like a=v/t?

OpenStudy (dan815):

power is the rate of work being done work=force*distance work/time = Power

OpenStudy (dan815):

work/time = force*distance/time but distance/time = speed so force * V = power 1100kg gives u 1100kg*9.81m/s^2 = approximately 11000 Newtons of Force

OpenStudy (dan815):

11000 * speed= 2200 watts

OpenStudy (karlaltr):

i think that makes sense thank you!

OpenStudy (dan815):

ok good

OpenStudy (dan815):

you can see the units match up too 11000 * speed= 2200 watts newtons * m/s = newtos *m/s watts=work/time = Force*distance/time = newtons * meters / sec